Evaluate 6^x if x =-2
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)
Answer by stanbon(75887)

Evaluate 6^x if x =-2
---
6^(-2) = 1/36
=================
Cheers,
Stan H.
=============